• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Higinbotham

    Log entry time 01:18:25 on October25,2001

    Entry number 74960

    keyword=Q1 Problem Continues

    The problem with Q1 continues. After talking with Cryo and the tech on call, it
    was clear that fixing this problem was probably going to take an access. Since we
    have many items that can be done tonight without the right arm, we will let MCC
    continue work on the beam and do our beam calibration runs. In the morning,
    when Cryo and the Hall A techs are here, we will fix the problem with Q1.